A cardinal flies east for 2.9 s in a horizontal plane for a distance of 22 m from a fence post a bush. It then flies north another 11 m to a bird feeder 1.5 s. Calculate average velocity of the cardin


1
Expert's answer
2021-02-09T03:46:16-0500

The average velocity equals:

V = (d1 + d2)/ (t1 + t2) = (22 m + 11 m) / (2.9 s + 1.5 s) = 7.5 m/s


Answer: 7.5 m/s